OpenStudy (vortish):

solve the equation (x-8)(3x+5)=0

OpenStudy (anonymous):

x=8 or -5/3

OpenStudy (anonymous):

just do x-8=0 and 3x+5=0

OpenStudy (anonymous):

4x-3=0 +3 +3 4x=3 divide by 4 on both sides x=3/4